OpenStudy (anonymous):

Find the slope of the line that passes through the following points (5,2) and (-3,-2)

OpenStudy (watchmath):

\(m=\frac{y_2-y_1}{x_2-x_1}=\frac{-2-2}{-3-5}=\frac{-4}{-8}=1/2.\)

OpenStudy (anonymous):

Thanks